OpenStudy (anonymous):

Write y =1/6 x + 12 in standard form using integers.

OpenStudy (anonymous):

The standard form is Ax + By = C -12 = 1/6 x - y 1/6 x - y = -12 x - 6y =-72

OpenStudy (zepp):

The stardard form should be Ax + By + C =0 So x - 6y + 72 = 0